# 2] organization=Observatoire des Sciences de l’Univers Nantes Atlantique (Osuna), addressline=CNRS UAR-3281, Nantes Université, country=France

### v2

@@ 5 Conclusions
We suggest the existence of an ordered structure within the E ring acting as a reflection grating with a millimetre-scale groove spacing. The structure, constrained between Saturn’s G-ring and Rhea’s orbit, likely consists of fresh ice particles supplied by Enceladus’ plumes. However, questions persist regarding the higher concentration of larger particles in the diffracting structure compared to those in the plumes. In addition, the stripe inclination of 43 to Saturn’s ring plane is unusual for particles within the E ring. Potential explanations may involve preferential ice crystal orientations, particle shapes, or the influence of E ring tendrils.
Recommendations for future research include filtering the VIMS cube to enhance the stripe contrast and reduce noise. This will aid in producing clearer spectral kymographs and deconstructing the contributions of the groove spacings and widths of the reflection grating.
@@
The origin of the diffraction effects can be further explored through diffraction analysis of ice crystal shapes, as well as simulations of crystal orientations and vertical tendril structures within the E ring.
@@
A survey of ISS and VIMS-IR data was conducted to find stripe features similar to those observed during flybys E13, E17, and E19. Expanding this survey to include VIMS-IR non-targeted flybys at high phase angles, primarily when Cassini is located between the G-ring and Rhea’s orbit, is advised. Investigating other instruments with similar near-infrared sensitivity and observation geometries would also be valuable.

### v3

@@ 5 Conclusions
We suggest the existence of an ordered structure within the E ring acting as a reflection grating with a millimetre-scale groove spacing. The structure, constrained between Saturn’s G ring and Rhea’s orbit, likely consists of fresh ice particles supplied by Enceladus’ plumes. However, questions persist regarding the higher concentration of larger particles in the diffracting structure compared to those in the plumes. In addition, the stripe inclination of 43 to Saturn’s ring plane is unusual for particles within the E ring. Potential explanations for the observed chromatic stripes include periodic enhancements in the E ring’s particle density—possibly caused by structures such as tendrils, wakes, or dust acoustic waves. Alternatively, periodic aggregates could also produce the required diffraction effects.
Recommendations for future research include filtering the VIMS cube to enhance the stripe contrast and reduce noise. This will aid in producing clearer spectral kymographs and deconstructing the contributions of the groove spacings and widths of the grating structure.
@@
The origin of the diffraction effects can be further investigated through diffraction analyses of ice crystal shapes and aggregates, as well as simulations of crystal orientations, vertical tendril structures, wakes, and dust acoustic waves within the E ring.
@@
A survey of ISS and VIMS-IR data was conducted to find stripe features similar to those observed during flybys E13, E17, and E19. Expanding this survey to include VIMS-IR non-targeted flybys at high phase angles, primarily when Cassini is located between the G ring and Rhea’s orbit, is advised. Investigating other instruments with similar near-infrared sensitivity and observation geometries would also be valuable.